# On modules whose dual is of finite Gorenstein dimension: The $k$-torsion, homological criteria, and applications in modules of differentials of order $n$

### v4

@@ 6. Gorenstein criteria and related questions
Let $R$ be a local ring. Suppose $\operatorname{depth}_{R}(R)=0$. By the Auslander-Bridger formula, all $R$-modules of finite Gorenstein dimension are totally reflexive. Moreover of for any $R$-module $M$ we have that the following implications hold:
$\operatorname{G-dim}_{R}(M^{\ast})<\infty$ $\Rightarrow\operatorname{G-dim}_{R}(\operatorname{Tr}_{R}(M))<\infty$
$\Rightarrow\operatorname{G-dim}_{R}(\operatorname{Tr}_{R}(M))=0$
$\Rightarrow\operatorname{G-dim}_{R}(M)=0.$
In particular, $R$ satisfies (GDUAL). Hence the existence of non-Gorenstein local rings of depth zero implies that Question 6.9 is false in general.
Motivated by condition $(\operatorname{GDUAL})$, we define a weaker condition than it, and prove that if $R$ is of positive depth and satisfies such condition, then $R$ is Gorenstein. This will show that Question 6.9 is positive for local rings of positive depth.
@@
Now, motivated by Corollary 5.14, we will discuss about a question that was addressed by Holanda and Miranda-Neto in [28]. The question is as follows:
@@
([28, Question 5.24]) Let $R$ be a Cohen-Macaulay local ring with a canonical module $\omega$. If $\operatorname{pd}_{R}(\omega^{\ast})<\infty$, then must $R$ be Gorenstein?
Firstly, we observe that the answer to Question 6.12 is positive and follows directly from the Corollary 5.14. However, it is worth highlighting that this question had already been answered affirmatively, albeit through a different proof, by Asgharzadeh in [4, Theorem 11.3]. Additionally, it is worth noting that long before Holanda and Miranda-Neto posed this question, the same result could be obtained directly from the Foxby equivalence (see [9, Theorem 3.4.11]).
We finish this section by proposing the following question, which is motivated by a celebrated result of Foxby [17] that asserts that the existence of a non-zero (finitely generated) $R$-module $M$ of projective dimension and injective dimension both finite implies that $R$ is Gorenstein.
@@
If there exists a (finitely generated) $R$-module $M$ such that $\operatorname{G-dim}_{R}(M^{\ast})<\infty$ and $\operatorname{id}_{R}(M)<\infty$, then is $R$ Gorenstein?
@@ 7. Preliminary facts on Kähler differentials
In this section, we establish our notation and, for the reader’s convenience, we recall some known results concerning Kähler differential modules and differential modules, both in the context of algebras and locally ringed spaces.
Affine derivation modules and Kähler fifferentials $n$-th order
Let $R$ be a $S$-algebra and $M$ be an $R$-module. Recall that a $S$-linear map $D:R\rightarrow M$ is said to be a derivation if for any two elements $x_{0},x_{1}\in R$, the following identity holds:
@@
A $S$-linear map $D:R\rightarrow M$ is said to be a $n$-th order derivation if for any $x_{0},\ldots,x_{n}\in R$, the following identity holds:
@@
Define the map $d_{n}^{R}:R\rightarrow\Omega^{(n)}_{R/S}$,  $a\mapsto(1\otimes a-a\otimes 1)+I^{n+1}$. This map is a derivation of order $n$, and its image generates $\Omega^{(n)}_{R/S}$ as an $R$-module (see [41, Chapter II-1]).
The $R$-module $\Omega^{(n)}_{R/S}$ is called the module of Kähler differentials of order $n$ of $R$ over $S$. The map $d_{n}^{R}$ is called the canonical derivation of $R$ in $\Omega^{(n)}_{R/S}$. It comes equipped with a universal derivation $d_{S/R}\in\text{Der}_{R}^{n}(S,\Omega^{(n)}_{S/R})$ with the property that composition with $d_{S/R}$ yields an isomorphism $\operatorname{Hom}_{R}(\Omega^{(n)}_{R/S},R)\cong\text{Der}_{S}^{n}(R)$ ([43, Proposition 1.6]). To see more properties regarding modules of derivations and Kähler differentials, we recommend [43, 51, 41].
@@
It is important to note that the differential module and Kähler differentials may not be finitely generated. However, they are finitely generated in certain cases. For instance: If $R$ is essentially of finite type over $S$. If $S=k$ is a field with a valuation and $R$ is an analytic $k$-algebra, meaning $R$ is module-finite over a convergent power series ring $k\{x_{1},\ldots,x_{n}\}$. If $S=k$ is a field, $(R,\mathfrak{m})$ is a complete local ring, and $R/\mathfrak{m}$ is a finite extension of $k$. For this reason, in this article, for each $n$, we will consider $\Omega^{(n)}_{R/S}$ to be finitely generated as an $R$-module. In particular, by the universal property, $\text{Der}^{n}_{S}(R)$ is also finitely generated as an $R$-module.
Next, we collect some definitions and properties that will play a central role in this paper.
@@
A ringed space $(X,\mathcal{O}_{X})$ consists of a topological space $X$ paired with a sheaf of rings $\mathcal{O}_{X}$, known as the structure sheaf of $X$. If $\mathcal{O}_{X,x}$ is a local ring for every $x\in X$, $(X,\mathcal{O}_{X})$ is termed a locally ringed space. Let $\mathfrak{m}_{x}$ denote the maximal ideal of $\mathcal{O}_{X,x}$; $k(x):=\mathcal{O}_{X,x}/\mathfrak{m}_{x}$ is referred to as the residue field of $X$ at $x$. We consider the following locally ringed spaces: locally Noetherian schemes (i.e., a scheme is locally Noetherian if it has a covering by open affine subsets $\operatorname{Spec}R_{i}$, where each $R_{i}$ is a Noetherian ring) and analytic spaces (the ringed space $(X,\mathcal{O}_{X})$ is called an analytic space if every point $x\in X$ has a neighborhood $V$ such that $(V,\mathcal{O}_{X}|_{V})$ is isomorphic to some analytic model space, where $\mathcal{O}_{X,x}\cong\frac{\mathcal{O}_{\mathbb{C}^{n},x}}{{\mathcal{I}}_{x}}$ for all $x\in X$, and $\frac{\mathcal{O}_{\mathbb{C}^{n},x}}{{\mathcal{I}}_{x}}$ is Noetherian and reduced; for further details, e.g., see [29], [23]).
@@ 7.1. Global Kähler differentials
Next, we will recall the globalized definition of the constructions given above. To do this, instead of gluing together the modules $\Omega^{(1)}_{S/R}$, we provide a global definition, and then point out that it reduces to the original definition given above. Before defining it, it is worth noting that the fiber product exists in our context (see, for example [10, Corollary 0.32] and [20, Theorem 4.18]). Let $\varphi:X\to Y$ be a morphism of space. This induces a morphism $\Delta:X\to X\times_{Y}X$, called the diagonal morphism. The map $\Delta$ comes with a map of sheaves $\Delta^{\#}:\mathcal{O}_{X\times_{Y}X}\to\Delta_{*}\mathcal{O}_{X}$ on $X\times_{Y}X$. Although really a map of sheaves of rings, we will regard it as a map of sheaves of $\mathcal{O}_{X\times_{Y}X}$-modules. Let $\mathcal{I}$ be its kernel, again regarded as an $\mathcal{O}_{X\times_{Y}X}$-module.
@@
The sheaf relative Kähler differential of $\mathcal{O}_{X}$-module is defined as:
@@
Note that if $Y=\operatorname{Spec}(k)$, where $k$ is a field, $\Omega_{X/Y}^{(n)}$ is usually denoted by $\Omega_{X}^{(n)}$. For further details and properties, see for example, [1, Section I.6.6, p. 34].
@@
Let $X$ and $Y$ be (locally) Noetherian spaces.
@@
    ([34, Proposition 6.1.20]). If $\varphi:X\to Y$ is locally of finite type, then $\Omega_{X/Y}^{(1)}$ is coherent. Additionally, by [20, Proposition 7.48], $\Omega_{X/Y}^{(n)}$ is also $\mathcal{O}_{X}$-coherent for all $n>1$.
@@
    ([34, Proposition 1.17]). If $\varphi:X\to Y$ is a morphism of spaces, then for $x\in X$ there is a canonical isomorphism of $\mathcal{O}_{X,x}$-modules
    $\left(\Omega_{X/Y}^{(1)}\right)_{x}=\Omega_{\mathcal{O}_{X,x}/\mathcal{O}_{Y,\varphi(x)}}^{(1)}.$
    Since the stalk commutes with exterior power, more generally, one obtains $(\Omega_{X/Y}^{(n)})_{x}\cong\Omega_{\mathcal{O}_{X,x}/\mathcal{O}_{Y,\varphi(x)}}^{(n)}=:\Omega_{(X,x)/(Y,\varphi(x)}^{(n)}$.
@@
    ([1, Corollaire (I6.5.5)]). There is also a definition of the sheaf of $Y$-derivations over $X$, denoted by $\mathcal{D}er_{Y}^{1}(X)$. Moreover, the universal property of $\Omega_{X/Y}^{(1)}$ comes down to saying that there is an isomorphism of $\mathcal{O}_{X}$-modules $\mathcal{D}er_{Y}^{1}(X)\cong\mathcal{H}om_{\mathcal{O}_{X}}\left(\Omega_{X/Y}^{(1)},\mathcal{O}_{X}\right).$ Naturally, is defined the sheaf of $n$-th order $Y$-derivations over $X$.
    $\mathcal{D}er_{Y}^{n}(X):=\mathcal{H}om_{\mathcal{O}_{X}}\left(\Omega_{X/Y}^{(n)},\mathcal{O}_{X}\right).$
@@
    [46, Lemma 28.20.2, Section 05P1] Let $X$ be a locally Noetherian scheme, and let $\mathcal{F}$ be a coherent $\mathcal{O}_{X}$-module. Then $\mathcal{F}$ is locally free if and only if for all $x\in X$, the stalk $\mathcal{F}_{x}$ is a free $\mathcal{O}_{X,x}$-module.
Based on Remarks 7.1 and Facts 7.4, the schemes considered in this section will be locally Noetherian, and the $\mathcal{O}_{X}$-module $\mathcal{D}er_{Y}^{n}(X)$ will also always be considered a non-zero coherent sheaf for each integer $n\geq 1$. Additionally, we assume that the sheaves have non-zero stalks, implying that the chosen elements belong to the support. For simplicity, the stalk of $\Omega_{(X,x)/(Y,\varphi(x))}^{(n)}$, as defined in Facts 7.4(2), will be denoted by $\Omega_{(X,x)/(Y,y)}^{(n)}$, where $y=\varphi(x)$.
@@
Let $X$ and $Y$ be locally Noetherian spaces, then the following properties hold:
@@
    $\mathcal{D}er_{Y}^{n}(X)$ is a $\mathcal{O}_{X}$-coherent.
@@
    for $x\in X$ there is a canonical isomorphism of $\mathcal{O}_{X,x}$-modules
    $(\mathcal{D}er_{Y}^{n}(X))_{x}\cong{\rm Der}^{n}_{\mathcal{O}_{Y,y}}(\mathcal{O}_{X,x}).$
    Thus, ${\rm Der}^{n}_{\mathcal{O}_{Y,y}}(\mathcal{O}_{X,x})$ will be denoted by ${\rm Der}^{n}_{{(Y,y)}}(X,x).$
@@
(i) From Fact 7.4(1), we establish the coherence of $\Omega_{X/Y}^{(n)}$. Subsequently, from that $\mathcal{D}er_{Y}^{n}(X)=\mathcal{H}om_{\mathcal{O}_{X}}\left(\Omega_{X/Y}^{(n)},\mathcal{O}_{X}\right)$ and by [34, Exercise 5.1.6(b)], we deduce the result.
@@
Since $\Omega_{X/Y}^{(n)}$ is $\mathcal{O}_{X}$-coherent (Facts 7.4(1)), so $\mathcal{H}om$ commutes with the stalk [20, Proposition 7.27]. Now from Facts 7.4(2)(3) from the universal property ( [27, Theorem 2.2.6] or [43, Proposition 1.6]) we get the last isomorphism. ∎
8. Applications on the $k$-torsion of the modules of differentials
Next, we study when the Kähler differential module $\Omega_{X/Y}^{(n)}$ of order $n$ is locally free, reflexive, $k$-torsion-free and $k$-syzygy, specially when $k=1,2$. The study of these properties in the modules of Kähler differential of order $n$ concerning the regularity questions of affine rings has a long history; when $n=1$ and $Y$ is a Spec of a field (see Kunz-Waldi’s book [31]). Next, we provide a summary given by Milher in [39], which is relevant for motivating the study of these concepts. Let $d:=\dim_{x}X=\dim(\mathcal{O}_{X,x})$ and $s:=\operatorname{codim}_{x}\operatorname{Sing}X$ (definition for ${\rm Sing}$ and $\operatorname{codim}$ see 8.6). Lipman (see also Suzuki [47]) showed that for an affine complete intersection germ over a field of characteristic zero, $\Omega^{(1)}_{X,x}$ is torsion-free if and only if $X$ is normal at $x\in X$, while $\Omega^{(1)}_{X,x}$ is reflexive if and only if $s>3$ (see [33]). Vetter proved in [49] that for a non-smooth reduced complete intersection singularity and for $1\leq n\leq d$, the module $\Omega^{(n)}_{X,x}$ is torsion-free (resp., reflexive) if $n<s$ (resp., $n<s-1$) and has torsion (resp., cotorsion) if $n=s$ (resp., if $n=\max\{1,s-1\}$). Lebelt strengthened Vetter’s result by showing that if $\Omega^{(n)}_{X,x}$ is torsion-free (resp., reflexive), then $n<s$ (resp., $n<s-1$) (see [32]). Greuel in [24] show that $\Omega^{(n)}_{X,x}$ has torsion for any $n>d$; (for any reduced complex analytic space, [16]). This fact implies that $(\Omega^{(n)}_{X,x})^{**}=0$ since $d>0$ and consequently results in $\text{cotor}\Omega^{(n)}_{X,x}=0$ for $n>d$. Also, new results on the connection between torsion and cotorsion and reflexivity of Kähler differential modules can be found in [22, 39]. Due to it and the results mentioned above, we obtain the following results:

### v5

@@ 6. Gorenstein criteria and related questions
If the local ring $R$ satisfies $\operatorname{(GDUAL)}$, then $R$ is not necessarily Gorenstein unless $\operatorname{depth}_{R}(R)>0$. Indeed, suppose $\operatorname{depth}_{R}(R)=0$. From the Auslander-Bridger formula, we can see that all $R$-modules of finite Gorenstein dimension are totally reflexive. Thus, any $R$-module whose dual has finite Gorenstein dimension, is totally reflexive (see Remark 2.4 and Facts 2.3(2)), and hence $R$ satisfies $\operatorname{(GDUAL)}$. Hence the existence of non-Gorenstein local rings of depth zero shows that $R$ is not necessarily Gorenstein. The case $0pt_{R}(R)>0$ is discussed in the next proposition.
Motivated by condition $(\operatorname{GDUAL})$, we define a weaker condition than it, and prove that if $R$ is of positive depth and satisfies such condition, then $R$ is Gorenstein.
@@
Now, motivated by Corollary 5.13, we will discuss about a question that was addressed by Holanda and Miranda-Neto in [VanishingOf(Co)homologyfrenessCriteriaAndTheAuslanderReitenConjectureforCohen-MacaulayBurch]. The question is as follows:
@@
([VanishingOf(Co)homologyfrenessCriteriaAndTheAuslanderReitenConjectureforCohen-MacaulayBurch, Question 5.24]) Let $R$ be a Cohen-Macaulay local ring with a canonical module $\omega$. If $\operatorname{pd}_{R}(\omega^{\ast})<\infty$, then must $R$ be Gorenstein?
Firstly, we observe that the answer to Question 6.10 is positive and follows directly from the Corollary 5.13. However, it is worth highlighting that this question had already been answered affirmatively, albeit through a different proof, by Asgharzadeh in [ReflexivityRevisited, Theorem 11.3]. Additionally, it is worth noting that long before Holanda and Miranda-Neto posed this question, the same result could be obtained directly from the Foxby equivalence (see [GorensteinDimensions, Theorem 3.4.11]).
We finish this section by proposing the following question, which is motivated by a celebrated result of Foxby [FoxbyIsomorphimsbetween] as well as a result of Holm [RingswithfiniteGorensteininjectivedimension]. The first (resp. second) asserts that the existence of a non-zero (finitely generated) $R$-module $M$ of projective dimension (resp. Gorenstein dimension) and injective dimension both finite implies that $R$ is Gorenstein.
@@
[…]
@@ 7. Preliminary facts on Kähler differentials
In this section, we introduce some notations and recall known results on Kähler and differential modules in the context of algebras and locally ringed spaces for convenience.
Affine derivation modules and Kähler differentials $n$-th order
Let $S$ be a (commutative) ring and $R$ be an $S$-algebra and $M$ be an $R$-module. Recall that an $S$-linear map $D:R\rightarrow M$ is said to be a derivation if for any two elements $x_{0},x_{1}\in R$, the following identity holds:
@@
An $S$-linear map $D:R\rightarrow M$ is said to be an $n$-th order derivation if for any $x_{0},\ldots,x_{n}\in R$, the following identity holds:
@@
Define the map $d_{n}^{R}:R\rightarrow\Omega^{(n)}_{R/S}$,  $a\mapsto(1\otimes a-a\otimes 1)+I^{n+1}$. This map is a derivation of order $n$, and its image generates $\Omega^{(n)}_{R/S}$ as an $R$-module (see [Nakai, Chapter II-1]).
The $R$-module $\Omega^{(n)}_{R/S}$ is called the module of Kähler differentials of order $n$ of $R$ over $S$. The map $d_{n}^{R}$ is called the canonical derivation of $R$ in $\Omega^{(n)}_{R/S}$. It comes equipped with a universal derivation $d_{S/R}\in\text{Der}_{R}^{n}\left(S,\Omega^{(n)}_{S/R}\right)$ with the property that composition with $d_{S/R}$ yields an isomorphism $\operatorname{Hom}_{R}\left(\Omega^{(n)}_{R/S},R\right)\cong\text{Der}_{S}^{n}(R)$ ([O, Proposition 1.6]). To see more properties regarding modules of derivations and Kähler differentials, we recommend [O, NKI, Nakai].
@@
It is important to note that the modules of derivations and the Kähler differentials may not be finitely generated. However, they are finitely generated in certain cases. For instance: If $R$ is essentially of finite type over $S$. If $S=k$ is a field with a valuation and $R$ is an analytic $k$-algebra, meaning $R$ is module-finite over a convergent power series ring $k\{x_{1},\ldots,x_{n}\}$. If $S=k$ is a field, $(R,\mathfrak{m})$ is a complete local ring, and $R/\mathfrak{m}$ is a finite extension of $k$. For this reason, in this article, for each $n$, we will consider $\Omega^{(n)}_{R/S}$ to be finitely generated as an $R$-module. In particular, by the universal property, $\text{Der}^{n}_{S}(R)$ is also finitely generated as an $R$-module.
We now present key definitions and properties that are central to this section.
@@
A scheme is a ringed topological space $(X,\mathcal{O}_{X})$ admitting an open covering $\{U_{i}\}_{i}$ such that $(U_{i},\mathcal{O}_{X}|_{U_{i}})$ is an affine scheme for every $i$. Therefore, a scheme is a locally ringed space. A scheme $(X,\mathcal{O}_{X})$ is said to be Noetherian if it is a finite union of affine open $\operatorname{Spec}(R_{i})$ such that $R_{i}$ is a (commutative) Noetherian ring for every $i$. We say that a scheme is locally Noetherian if every point has a Noetherian open neighborhood.
We can see that if $(X,\mathcal{O}_{X})$ is a locally Noetherian scheme, then the local rings $\mathcal{O}_{X,x}$ are Noetherian rings. Often we simply write $X$ instead of $(X,\mathcal{O}_{X})$.
@@ 7.1. Global Kähler differentials
Next, we recall the global definition of the constructions given above. To do this, instead of gluing together the modules $\Omega^{(1)}_{S/R}$, we provide a global definition, and then point out that it reduces to the original definition given above. Before defining it, it is worth noting that the fiber product exists in our context (see [GW, Theorem 4.18]). Let $\varphi:X\to Y$ be a morphism of schemes. This induces a morphism $\Delta:X\to X\times_{Y}X$, called the diagonal morphism. The map $\Delta$ comes with a map of sheaves $\Delta^{\#}:\mathcal{O}_{X\times_{Y}X}\to\Delta_{*}\mathcal{O}_{X}$ on $X\times_{Y}X$. Although really a map of sheaves of rings, we will regard it as a map of sheaves of $\mathcal{O}_{X\times_{Y}X}$-modules. Let $\mathcal{I}$ be its kernel, again regarded as an $\mathcal{O}_{X\times_{Y}X}$-module.
@@
The sheaf relative Kähler differential of $\mathcal{O}_{X}$-modules is defined as:
@@
Note that if $Y=\operatorname{Spec}(k)$, where $k$ is a field, $\Omega_{X/Y}^{(n)}$ is usually denoted by $\Omega_{X}^{(n)}$. For further details and properties, see for example, [Gro1, Section I.6.6, p. 34].
@@
Let $\varphi:X\to Y$ be a morphism of locally Noetherian schemes.
@@
    ([Qing, Proposition 6.1.20] or [Gro1, Corollaire (I6.4.22)]). If $\varphi:X\to Y$ is locally of finite type, then $\Omega_{X/Y}^{(1)}$ is coherent. Additionally, by [GW, Proposition 7.48], $\Omega_{X/Y}^{(n)}$ is also a coherent $\mathcal{O}_{X}$-module for all $n>1$.
@@
    ([Qing, Proposition 1.17, ch. 6]). For each $x\in X$, there is a canonical isomorphism of $\mathcal{O}_{X,x}$-modules (Ω_X/Y^(1))_x=Ω_O_X,x/O_Y,φ(x)^(1). Since the stalk commutes with exterior power, more generally, one obtains $\left(\Omega_{X/Y}^{(n)}\right)_{x}\cong\Omega_{\mathcal{O}_{X,x}/\mathcal{O}_{Y,\varphi(x)}}^{(n)}=:\Omega_{(X,x)/(Y,\varphi(x))}^{(n)}$.
@@
    ([Gro1, Corollaire (I6.5.5)]). There is also a definition of the sheaf of $Y$-derivations over $X$, denoted by $\mathcal{D}er_{Y}^{1}(X)$. Moreover, the universal property of $\Omega_{X/Y}^{(1)}$ comes down to saying that there is an isomorphism of $\mathcal{O}_{X}$-modules $\mathcal{D}er_{Y}^{1}(X)\cong\mathcal{H}om_{\mathcal{O}_{X}}\left(\Omega_{X/Y}^{(1)},\mathcal{O}_{X}\right).$ Naturally, is defined the sheaf of $n$-th order $Y$-derivations over $X$. Der_Y^n(X):=Hom_O_X(Ω_X/Y^(n),O_X).
@@
    ([stacks-project, Lemma 28.20.2, Tag 05P1]). Let $X$ be a locally Noetherian scheme, and let $\mathcal{F}$ be a coherent $\mathcal{O}_{X}$-module. Then $\mathcal{F}$ is locally free if and only if for all $x\in X$, the stalk $\mathcal{F}_{x}$ is a free $\mathcal{O}_{X,x}$-module.
We provide an elementary lemma that may be familiar, but we have not found it in the literature.
@@
Let $X$ and $Y$ be locally Noetherian schemes, and let $\varphi:X\to Y$ be a morphism locally of finite type. Then, the following properties hold:
@@
    $\mathcal{D}er_{Y}^{n}(X)$ is a coherent $\mathcal{O}_{X}$-module.
@@
    For $x\in X$ there is a canonical isomorphism of $\mathcal{O}_{X,x}$-modules (Der_Y^n(X))_x≅Der^n_O_Y,φ(x)(O_X,x).
@@
(i) From Facts 7.4(1), we establish the coherence of $\Omega_{X/Y}^{(n)}$. Subsequently, from the equality $\mathcal{D}er_{Y}^{n}(X)=\mathcal{H}om_{\mathcal{O}_{X}}\left(\Omega_{X/Y}^{(n)},\mathcal{O}_{X}\right)$ and [Qing, Exercise 5.1.6(b)], we deduce the result.
@@
Since $\Omega_{X/Y}^{(n)}$ is coherent $\mathcal{O}_{X}$-module (Facts 7.4(1)), so $\mathcal{H}om$ commutes with taking stalks ([GW, Proposition 7.27]). Now from Facts 7.4(2)(3) from the universal property ( [HS, Theorem 2.2.6] or [O, Proposition 1.6]) we get the last isomorphism. ∎
Based on Remarks 7.1, Facts 7.4, and Lemma 7.5, the schemes considered will be locally Noetherian, and the sheaves $\mathcal{D}er_{Y}^{n}(X)$ and $\Omega_{X/Y}^{(n)}$ will always be non-zero coherent sheaves. Furthermore, we assume that these sheaves have non-zero stalks, implying that the chosen elements lie in the support. For simplicity, the stalks of $\Omega_{(X,x)/(Y,\varphi(x))}^{(n)}$ and ${\rm Der}^{n}_{\mathcal{O}_{Y,\varphi(x)}}(\mathcal{O}_{X,x})$ will be denoted by $\Omega_{(X,x)/(Y,y)}^{(n)}$ and ${\rm Der}^{n}_{(Y,y)}(X,x)$, respectively, where $y=\varphi(x)$.
8. Applications
8.1. Applications on the $k$-torsion of the modules of differentials
Next, we study the conditions under which the Kähler differential module $\Omega_{X/Y}^{(n)}$ over schemes is reflexive, $k$-torsionfree, and $k$-syzygy, with particular attention to the cases where $k=1$ and $k=2$. The investigation of these properties in Kähler differential modules of order $n$ concerning regularity questions of affine rings has a long history in the literature, addressed by various authors, especially in the context of algebraic varieties and analytic spaces, particularly when $Y$ is the spectrum of field, see, for example, [Kunz, Milher, graf, Lipman, Suzuki, Vetter, Lebelt, Greuel]). Recently, Graf and Milher in [graf, Milher] gave new results on the connection between torsion, cotorsion, and reflexivity of the Kähler differential modules of order $n$ in complex analytic varieties and algebraic varieties. For a summary of the existing results, see [Milher, p. 133]. Motivated by these results, we provide the following results in the context of schemes.